Casper Landsten once again has ​$0.9million​ (or its Swiss franc​ equivalent) to invest for three months. He now faces the following rates. Should he enter into a covered interest arbitrage​ (CIA) investment?

Arbitrage funds available $900,000

Spot exchange rate (sFr/$) 1.3394

3-month forward rate (sFr/$) 1.3287

US Dollar Int. rate 4.752%

Swiss Franc annual Int rate 3.627%

The CIA Profit potential is ____%, which tells Casper Landsten he should borrower ______(Swiss Francs or US Dollars) and invest in the ____(higher or lower) yielding currency, the ____(Swiss Franc or US Dollar) and then sell the ____ (Swiss Franc or US Dollar) principal and interest forward 3 months locking in a CIA profit. (Round to three decimals and select from the drop downs. - Drop Down selects are in parathesis
Answer #1

Spot exchange rate : SFr /$=1.3394

Future Value of 1$=1*(1+(0.04752/4))=1.01188

Future Value of 1.3394 SFr=1.3394*(1+(0.03627/4))=1.351545

Fair spot rate after 3 months :

1.351545SFr=1.01188 $

1$=1.351545/1.01188=1.335677SFr

Forward Rate:

1$=1.3287SFr

Forward rate SFr is costlier

Hence SFr needs to be sold in forward for arbitrage

Arbitrage Strategy:

Borrow US dollar at interest rate=4.752%

Convert to SFr at current spot rate:: SFr /$=1.3394

Invest SFr at interest rate=3.627%

Sell SFr and convert at forward rate:

1$=1.3287SFr

CALCULATION OF PROFIT POTENTIAL:

Amount of SFr received at current exchange rate:

$0.9 million*1.3394SFr=1.20546 SFr

Future Value of 1.20546 million SFr=1.20546*(1+(0.03627/4))=1.216391 Million SFr

Amount of $ received after conversion at Forward Rate=(1/1.3287)*1.216391 million=0.915475 Million $

Profit Potential =(0.915475/0.9)-1=0.017194=1.719%(rounding to two decimal)

The CIA Profit potential is 1.719%, which tells Casper Landsten he should borrower US Dollars and invest in the Lower yielding currency, the Swiss Franc and then sell the Swiss Franc principal and interest forward 3 months locking in a CIA profit. (Round to three decimals and select from the drop downs.